# Exact solution for the current generated by dissipative master equation in single-band tight-binding systems subject to time-dependent uniform electric fields

### v1

@@ I Introduction
Modelling the optical response of solid state materials is a cornerstone of modern optoelectronics. New devices are often based on desirable optical properties that can be engineered in synthetic materials [1]. This requires a fine understanding of the physics underlying the light-matter interactions in these systems.
@@
The investigation of optical properties in solid state systems amounts to determining the electrical currents generated by the electromagnetic field of an incomming light. For a long time, available light sources were weak, such that the current response of most materials could be described within the framework of linear response. With the advent of lasers [2], however, hitherto ignored nonlinear currents became a central feature of a new class of physical phenomena. Since then, strong-field phenomena in quantum systems became an active field of research [3, 4, 5, 6, 7, 8, 9]. A hallmark of nonlinear response is that in such phenomena, the system responds at frequencies that are multiples of the fundamental driving frequency. Second and third harmonic generation [10, 11], Monochromatic wave mixing and high-harmonic generation[9, 12] are examples of such nonlinear effects.
@@
When the driving field becomes strong, environmental dissipation effects play a vital role in determining the out-of-equilibrium state of the system.This problem has a long history of research, which has lead to the development of several frameworks to include dissipative mechanisms in quantum theories [13, 14, 15, 16, 17, 18, 19, 20, 21]. One widely used theoretical tool to formulate the dynamics of open systems is the Keldysh formalism [22], which utilises the nonequilibrium Green’s function (NEGF) [3, 23, 24, 25, 26, 27] as a fundamental quantity rather than the wave function. The NEGF formalism is ideally suited for the development of perturbative and self-consistent approximations. An alternative class of methods to describe open-quantum systems are density matrix equations of motions, typically referred to as Quantum Master Equations (QMEs)[28, 29]. Such equations can sometimes be reduced to the so called Lindblad equation, which describe Markovian evolution [28]. The simplest, but still useful, description of dissipation is provided by the Relaxation Time Approximation (RTA), which assumes that in the absence of drive the system relaxes to a known equilibrium state, in a time-scale determined by a relaxation rate[30, 31, 32, 33, 34]. The RTA has been widely used in conjunction with the Boltzmann equation [35, 36] and the semiconductor Bloch equations, on top of which perturbation theory schemes have been developed[37, 38, 39, 40, 41]. Despite its phenomenological nature, which can bring about unphysical behaviour in multiband systems [42], the RTA is suficient to capture a broad range of rich phenomena. Examples of these are the generation of direct
@@
An example of nonlinear optical effect is the production of direct currents driven by optical excitation, a class of responses that is collectively known as photogalvanic effects (PGEs) [43, 38, 44, 45, 46, 47]. This class of responses can be predicted with perturbation theory (on the electromagnetic field) within RTA [41], but the direct currents show anomalous dependencies on the relaxation time [48, 43, 47]. In particular, they diverge in the collisionless limit, which questions the validity of standard perturbation theory. Nonperturbative calculations with Floquet-Keldysh formalism instead predict that these currents converge to a finite value in this limit [49]. Other than this, direct comparisons between standard perturbative and nonperturbative methods, probing the validty of perturbation theory, are lacking.
@@
The existence of exact analytical results, even when obtained for some simplified model, is invaluable in physics, for their rarity and since they enable comparison with both experimental results and theoretical methods obtained via approximate and/or numerical methods. This is also true in the study of solid-state optoelectronic systems, due to the challenges present by treating both strong fields and dissipation in an adequate way. Although analytical solutions for certain driven dissipative systems have been derived, they typically apply only to the stationary regime under a constant external field [50, 51, 52]. In this paper, we present the exact solution for the time-resolved current response of a electronic system described by a single-band tight-binding system, subject to a uniform electric field with arbitrary time-dependency, described within a QME under the RTA. The paper is organized as follows: in Section II we detail the theoretical framework and model used. We outline the steps taken to obtain the solution for the current and deduce the analytical exact expressions for the current’s time evolution in two limiting cases: a monochromatic field and a constant field, both turned-on at an initial time $t=0$. In Section III we obtain the analytical expression for the strong-field response of the system, discussing both: (i) the effect of dissipation on High Harmonic Generation (HHG), and discuss the emergence of even harmonics in centrosymmetric systems; and (ii) the steady-state DC current obtained due to a constant electric field, where a generalization for the Esaki-Tsu equation is presented, valid for any single-band tight-binding model and arbitrary spatial dimensionality. In Section IV, we compare the perturbative results with the exact ones. Finally, we state our conclusions in Section V.
@@ II Model and analytical solution
where $H(t)$ is the time-dependent Hamiltonian for independent electrons, and $\left.\frac{\partial\rho\left(t\right)}{\partial t}\right|_{\text{scat}}$ includes the scattering induced by electron-electron, electron-phonon and electron-impurity interactions [53]. This type of equation is also referred to as a semiconductor Bloch equation. In principle, this type of equationcan be obtained from well-established treatments of many-particle systems, such as the many-particle density matrix framework [54, 55] or the Keldysh Green function method [56]. The simplest approximation to the scattering term is the so-called Relaxation Time Approximation (RTA), in which Eq. (1) reduces to[30, 31, 32, 33, 34]
@@
where $\beta$ is the inverse temperature and $\mu$ is the chemical potential. In the long time limit, we can have a zero external electric field, but a non-zero vector potential. The equilibrium density matrix, $\rho^{\text{eq}}(t)$, appearing in the relaxation term of Eq. (2) should take this into account by setting $\rho_{\mathbf{k}}^{\text{eq}}(t)=\rho_{\mathbf{k}+e\mathbf{A}\left(t\right)/\hbar}^{\text{eq}}$ , where we shift the Bloch momentum of the unperturbed equilibrium distribution by $e\mathbf{A}\left(t\right)$ [37]. This shift guarantees that $\lim_{t\rightarrow\infty}\rho_{\mathbf{k}}(t)=\rho_{\mathbf{k}+e\mathbf{A}\left(t\right)/\hbar}^{\text{eq}}$ is solution of Eq. (2) provided $\lim_{t\rightarrow\infty}\mathbf{E}(t)=\mathbf{0}$.
@@ II.1 General solution
where we used the fact that $de^{i\mathbf{k}\left(t\right)\cdot\mathbf{l}}/dt=\frac{d\mathbf{k}\left(t\right)}{dt}\cdot\mathbf{l}e^{i\mathbf{k}\left(t\right)\cdot\mathbf{l}}$ and that $\frac{d\mathbf{k}\left(t\right)}{dt}=\frac{e}{\hbar}\frac{d\mathbf{A}\left(t\right)}{dt}$ in the LHS through partial differentiation. Note that the equation contains the quantity $C_{\mathbf{l}}\equiv\int_{FBZ}d\mathbf{k}\rho_{\mathbf{k}\left(t\right)}^{\text{eq}}e^{i\mathbf{k}\left(t\right)\cdot\mathbf{l}}=\int_{FBZ}d\mathbf{k}\rho_{\mathbf{k}}^{\text{eq}}e^{i\mathbf{k}\cdot\mathbf{l}}$ which is constant in time. The fact that the quantities $\tilde{\mathcal{J}}_{\mathbf{l}}\left(t\right)$ are proportional to a sum of exponentials that contain all time dependency allows a set of decoupled, ordinary and linear differential equations for $\tilde{\mathcal{J}}_{\mathbf{l}}\left(t\right)$. The solution to these equations is:
@@ II.2 Monochromatic field
Where we have defined $J_{n}^{\mathbf{l}}\equiv J_{n}\left(\left|\omega_{B}^{\mathbf{l}}/\Omega\right|\right)$ to simplify notation, with $\omega_{B}^{\mathbf{l}}=\frac{e}{\hbar}El\cos\phi_{\mathbf{l}}$, $l\equiv\left|\mathbf{l}\right|$, $\cos\phi_{\mathbf{l}}\equiv\mathbf{l}\cdot\mathbf{A}\left(t\right)/\left(lA\left(t\right)\right)$ and $A\left(t\right)\equiv\left|\mathbf{A}\left(t\right)\right|$. Each contribution to the current $\mathcal{J}_{\mathbf{l}}\left(t\right)$ is composed of a persistent oscillating term and a transient term, which depends on the sudden switching of the electric field, that decays exponentially in time as $e^{-\gamma t}$. It is worth noting at this point that the solution displays inversion symmetry, since inverting the electric field also results in the complete inversion of the current (this is made evident by simply performing the change $\phi\rightarrow\phi+\pi$). This makes sense, since a single-band system always displays inversion symmetry. The Taylor expansion of the Bessel functions in the electric field also allows us to gauge this. Because they have a very particular form of expansion $J_{\alpha}\left(z\right)=\sum_{n=0}^{\infty}C_{n}z^{2n+\alpha}$, it is not hard to see from the coefficients in Eq. 14 that only odd powers of the electric field amplitude appear, which means that the current is inverted if the electric field is inverted.
@@
Notice that now, in the clean limit $\gamma\rightarrow 0$, the second and third terms of Eq. 13 do not decay and actually dominate the response since the first term vanishes in this limit. It is evident from Eq. 15 that one can generate even harmonics for $\phi\neq 0$. Although the emergence of even harmonics is usually thought to be prohibited in centrosymmetric systems, this effect is solely due to the sudden switching of the electric field. If the switching is rather done adiabatically, the even harmonic contributions vanish (see Appendix VI.1). By introducing dissipation, the system eventually loses memory of the electric field’s initial condition and the even harmonics vanish. Furthermore, for $\phi=0$, the response is constituted solely of odd harmonics of amplitude $\chi_{2j+1}^{\gamma=0,\mathbf{l},1}=2J_{2j+1}^{\mathbf{l}}$ , as reported in Ref. [57].
@@ II.3 Constant field
which as we will see is a generalization of the Esaki-Tsu relation [58]. The physical meaning of this is easier to unveil when considering a one dimensional tight binding chain with nearest neighbour hoppings. In that case we only have contribution from neighbours located at $l=1$ and the frequency $\omega_{B}^{\mathbf{l}}=\frac{e}{\hbar}E$ reduces to the Bloch frequency. The transient time is therefore composed of decaying Bloch Oscillations (BOs). These decaying BOs have also been predicted to occur in two-terminal mesoscopic setups subject to strong electric fields [59], where the BOs decay is caused by the coupling of the mesoscopic system to external leads. If no dissipation was present, $\gamma=0$, then Eq. 17 would give a total current of $J\left(t\right)=-2t_{1}C\sin\left(\omega_{B}t\right)$, which are the expected persistent BOs predicted for the one-dimensional tight binding chain [60, 61]. The constant, long-time current for the chain is $J^{\text{s}}=-2Ct_{1}\frac{E/\gamma}{\left(\frac{\hbar}{e}\right)^{2}+\left(E/\gamma\right)^{2}},$ which is the Esaki-Tsu relation that has been already obtained for one-dimensional nearest-neighbour tight binding lattices [51, 50, 52]. The result in Eq. 18 is the generalization of the Esaki-Tsu formula for an arbitrary single-band tight-binding system.
@@ III Results
The electric field is considered to lay along the direction of the first neighbor which lies in the x axis and we also measure the current along this direction. The presence of an additional hopping gives rise to competing scales that generate interesting phenomena [57].
@@ III.1 Monochromatic field
In Fig. 2 we plot the ratio between the odd harmonic Fourier amplitudes and the first-harmonic amplitude $\left|\Lambda_{j}^{\mathbf{l}}\right|/\left|\Lambda_{1}^{\mathbf{l}}\right|$ of a single neighbor for a set of dissipation values $\gamma/\Omega$ and for $\omega_{B}^{\mathbf{l}}/\Omega=3,12.2$ and $32$. A first thing to notice is how the amplitudes start dying out exponentially for harmonics of order superior than $\omega_{B}^{\mathbf{l}}$. By analysing Eq.(14), one can understand this behaviour as originating from the properties of Bessel functions. At a fixed argument $z$, as a function of the order $n$, $J_{n}\left(z\right)$ decays exponentially as $J_{n}\left(z\right)\sim e^{-n\ln\left(z^{-1}\right)}$ for large order $n$ [62]. Another interesting result is how for strong dissipation the higher harmonics are suppressed, when compared to the first harmonic. This result is in agreement with results from Ref. [63], where it is interpreted as being a consequence of the suppression of a particle’s longer trajectories in real-space via dissipation, which are responsible for hi˛gh frequency contributions. However, our results indicate that the suppression of higher harmonics with dissipation is not monotonous with increased dissipation. As shown in Fig. (2)b), intermediate harmonics can have their amplitude increased with dissipation.
@@
are much smaller than the odd ones. In Fig. 4, we plot the transient and long times Fourier spectra for a system with $\omega_{B}^{\mathbf{l}}/\Omega=1.54$, $\Omega=0.15$ and $\gamma=0.003$ (Fig. 4b) and $\omega_{B}^{\mathbf{l}}/\Omega=2.5$, $\Omega=0.13$ and $\gamma=0.004$ (Fig. 4c). The difference between the transient and long times spectra is the aquisition time of the signal, which is kept close or far away from the regime where the transient current dominates, respectively. We can detect a small DC current and the second harmonic in the transient spectrum in Fig. 4 b) and c) respectively, with amplitudes of roughly one and two orders of magnitude smaller than that of the first harmonic. As the aquisition time increases, the oscillatory long time state has more significant contribution and the even harmonics decrease with respect to the first harmonic.
@@
Figure 4: a) Time-evolution of current for $\omega_{B}^{\mathbf{l}}/\Omega=1.54$, $\Omega=0.15$ and $\gamma=0.003$. The black dashed line delimits the aquisition time of for the transient regime spectrum. b) and c) Transient and long times Fourier spectra for b) $\omega_{B}^{\mathbf{l}}/\Omega=1.54$, $\Omega=0.15$ and $\gamma=0.003$ and c) $\omega_{B}^{\mathbf{l}}/\Omega=2.5$, $\Omega=0.13$ and $\gamma=0.004$ in logarithmic scale. The aquisition times for the transient and long times spectra were b) $353$ and $10000$ c) $2000$ and $42400$, respectively. A small DC current and the second harmonic are visible in the transient spectrum in b) and c) respectively, with amplitudes of roughly one and two orders of magnitude smaller than that of the first harmonic. As the aquisition time increases, the oscillatory long time state has a more significant contribution and the even harmonics decrease with respect to the first harmonic.

### v2

# Exact nonlinear dynamics of single band metallic systems with dissipation: From optical to dc currents

@@ I Introduction
Modelling the optical response of solid state materials is a cornerstone of modern optoelectronics. New devices are often based on desirable optical properties that can be engineered in synthetic materials [zhang_review_2021]. This requires a fine understanding of the physics underlying the light-matter interactions in these systems.
@@
The investigation of optical properties in solid state systems amounts to determining the electrical currents generated by the electromagnetic field of an incomming light. For a long time, available light sources were weak, such that the current response of most materials could be described within the framework of linear response. With the advent of lasers [strickland_compression_1985], however, hitherto ignored nonlinear currents became a central feature of a new class of physical phenomena. Since then, strong-field phenomena in quantum systems became an active field of research [aoki_nonequilibrium_2014, oka_floquet_2019, eckardt_colloquium_2017, basov_towards_2017, de_la_torre_colloquium_2021, morimoto_geometric_2023, silva_high_2019]. A hallmark of nonlinear response is that in such phenomena, the system responds at frequencies that are multiples of the fundamental driving frequency. Second and third harmonic generation [mund_high-resolution_2018, kumar_third_2013], Monochromatic wave mixing and high-harmonic generation[silva_high_2019, ghimire_high-harmonic_2019] are examples of such nonlinear effects.
@@
When the driving field becomes strong, environmental dissipation effects play a vital role in determining the out-of-equilibrium state of the system.This problem has a long history of research, which has lead to the development of several frameworks to include dissipative mechanisms in quantum theories [weiss_quantum_2012, barreiro_transport_2009, vandecasteele_current-voltage_2010, fang_high-field_2011, li_nonequilibrium_2018, okamoto_nonequilibrium_2007, sugimoto_field-induced_2008, heidrich-meisner_nonequilibrium_2010, ikeda_nonequilibrium_2021]. One widely used theoretical tool to formulate the dynamics of open systems is the Keldysh formalism [sadovskii_diagram_2023], which utilises the nonequilibrium Green’s function (NEGF) [aoki_nonequilibrium_2014, blandin_localized_1976, jauho_time-dependent_1994, onoda_theory_2006, han_nonequilibrium_2018, tsuji_correlated_2008] as a fundamental quantity rather than the wave function. The NEGF formalism is ideally suited for the development of perturbative and self-consistent approximations. An alternative class of methods to describe open-quantum systems are density matrix equations of motions, typically referred to as Quantum Master Equations (QMEs)[breuer_theory_2002, alicki_quantum_2007]. Such equations can sometimes be reduced to the so called Lindblad equation, which describe Markovian evolution [breuer_theory_2002]. The simplest, but still useful, description of dissipation is provided by the Relaxation Time Approximation (RTA), which assumes that in the absence of drive the system relaxes to a known equilibrium state, in a time-scale determined by a relaxation rate[sato_nonlinear_2021, sato_light-induced_2019, nuske_floquet_2020, boyd_nonlinear_2008, yamada_density_2014]. The RTA has been widely used in conjunction with the Boltzmann equation [peres_optical_2014, mikhailov_non-linear_2007] and the semiconductor Bloch equations, on top of which perturbation theory schemes have been developed[passos_nonlinear_2018, passos_nonlinear_2021, ventura_gauge_2017, sipe_nonlinear_1993, sipe_second-order_2000]. Despite its phenomenological nature, which can bring about unphysical behaviour in multiband systems [terada_limitations_2024], the RTA is suficient to capture a broad range of rich phenomena. Examples of these are the generation of direct
@@
An example of nonlinear optical effect is the production of direct currents driven by optical excitation, a class of responses that is collectively known as photogalvanic effects (PGEs) [ventura_second_2020, passos_nonlinear_2021, sodemann_quantum_2015, chan_photocurrents_2017, de_juan_quantized_2017, cheng_intraband_2019]. This class of responses can be predicted with perturbation theory (on the electromagnetic field) within RTA [sipe_second-order_2000], but the direct currents show anomalous dependencies on the relaxation time [fregoso_jerk_2018, ventura_second_2020, cheng_intraband_2019]. In particular, they diverge in the collisionless limit, which questions the validity of standard perturbation theory. Nonperturbative calculations with Floquet-Keldysh formalism instead predict that these currents converge to a finite value in this limit [matsyshyn_rabi_2021]. Other than this, direct comparisons between standard perturbative and nonperturbative methods, probing the validty of perturbation theory, are lacking.
@@
The existence of exact analytical results, even when obtained for some simplified model, is invaluable in physics, for their rarity and since they enable comparison with both experimental results and theoretical methods obtained via approximate and/or numerical methods. This is also true in the study of solid-state optoelectronic systems, due to the challenges present by treating both strong fields and dissipation in an adequate way. Although analytical solutions for certain driven dissipative systems have been derived, they typically apply only to the stationary regime under a constant external field [han_solution_2013, kolovskyAtomicCurrentOptical2008, minot_quantum_2004]. In this paper, we present the exact solution for the time-resolved current response of a electronic system described by a single-band tight-binding system, subject to a uniform electric field with arbitrary time-dependency, described within a QME under the RTA. The paper is organized as follows: in Section II we detail the theoretical framework and model used. We outline the steps taken to obtain the solution for the current and deduce the analytical exact expressions for the current’s time evolution in two limiting cases: a monochromatic field and a constant field, both turned-on at an initial time $t=0$. In Section III we obtain the analytical expression for the strong-field response of the system, discussing both: (i) the effect of dissipation on High Harmonic Generation (HHG), and discuss the emergence of even harmonics in centrosymmetric systems; and (ii) the steady-state DC current obtained due to a constant electric field, where a generalization for the Esaki-Tsu equation is presented, valid for any single-band tight-binding model and arbitrary spatial dimensionality. In Section IV, we compare the perturbative results with the exact ones. Finally, we state our conclusions in Section V.
@@ II Model and analytical solution
where $H(t)$ is the time-dependent Hamiltonian for independent electrons, and $\left.\frac{\partial\rho\left(t\right)}{\partial t}\right|_{\text{scat}}$ includes the scattering induced by electron-electron, electron-phonon and electron-impurity interactions [cheng_third-order_2015]. This type of equation is also referred to as a semiconductor Bloch equation. In principle, this type of equationcan be obtained from well-established treatments of many-particle systems, such as the many-particle density matrix framework [haug_quantum_2009, malic_microscopic_2011] or the Keldysh Green function method [haug_quantum_2008]. The simplest approximation to the scattering term is the so-called Relaxation Time Approximation (RTA), in which Eq. (1) reduces to[sato_nonlinear_2021, sato_light-induced_2019, nuske_floquet_2020, boyd_nonlinear_2008, yamada_density_2014]
@@
where $\beta$ is the inverse temperature and $\mu$ is the chemical potential. In the long time limit, we can have a zero external electric field, but a non-zero vector potential. The equilibrium density matrix, $\rho^{\text{eq}}(t)$, appearing in the relaxation term of Eq. (2) should take this into account by setting $\rho_{\mathbf{k}}^{\text{eq}}(t)=\rho_{\mathbf{k}+e\mathbf{A}\left(t\right)/\hbar}^{\text{eq}}$ , where we shift the Bloch momentum of the unperturbed equilibrium distribution by $e\mathbf{A}\left(t\right)$ [passos_nonlinear_2018]. This shift guarantees that $\lim_{t\rightarrow\infty}\rho_{\mathbf{k}}(t)=\rho_{\mathbf{k}+e\mathbf{A}\left(t\right)/\hbar}^{\text{eq}}$ is solution of Eq. (2) provided $\lim_{t\rightarrow\infty}\mathbf{E}(t)=\mathbf{0}$.
@@ II.1 General solution
where we used the fact that $de^{i\mathbf{k}\left(t\right)\cdot\mathbf{l}}/dt=\frac{d\mathbf{k}\left(t\right)}{dt}\cdot\mathbf{l}e^{i\mathbf{k}\left(t\right)\cdot\mathbf{l}}$ and that $\frac{d\mathbf{k}\left(t\right)}{dt}=\frac{e}{\hbar}\frac{d\mathbf{A}\left(t\right)}{dt}$ in the LHS through partial differentiation. Note that the equation contains the quantity $C_{\mathbf{l}}\equiv\frac{V_{d}}{\left(2\pi\right)^{d}}\int_{FBZ}d\mathbf{k}\rho_{\mathbf{k}\left(t\right)}^{\text{eq}}e^{i\mathbf{k}\left(t\right)\cdot\mathbf{l}}=\frac{V_{d}}{\left(2\pi\right)^{d}}\int_{FBZ}d\mathbf{k}\rho_{\mathbf{k}}^{\text{eq}}e^{i\mathbf{k}\cdot\mathbf{l}}$ which is constant in time, where $V_{d}$ is the volume of the unit cell and $d$ is the dimension of the system. The fact that the quantities $\tilde{\mathcal{J}}_{\mathbf{l}}\left(t\right)$ are proportional to a sum of exponentials that contain all time dependency allows a set of decoupled, ordinary and linear differential equations for $\tilde{\mathcal{J}}_{\mathbf{l}}\left(t\right)$. The solution to these equations is:
@@ II.2 Monochromatic field
Where we have defined $J_{n}^{\mathbf{l}}\equiv J_{n}\left(\left|\omega_{B}^{\mathbf{l}}/\Omega\right|\right)$ to simplify notation, with $\omega_{B}^{\mathbf{l}}=\frac{e}{\hbar}El\cos\phi_{\mathbf{l}}$, $l\equiv\left|\mathbf{l}\right|$, $\cos\phi_{\mathbf{l}}\equiv\mathbf{l}\cdot\mathbf{A}\left(t\right)/\left(lA\left(t\right)\right)$ and $A\left(t\right)\equiv\left|\mathbf{A}\left(t\right)\right|$. Each contribution to the current $\mathcal{J}_{\mathbf{l}}\left(t\right)$ is composed of a persistent oscillating term and a transient term, which depends on the sudden switching of the electric field, that decays exponentially in time as $e^{-\gamma t}$. It is worth noting at this point that the solution displays inversion symmetry, since inverting the electric field also results in the complete inversion of the current (this is made evident by simply performing the change $\phi\rightarrow\phi+\pi$). This makes sense, since a single-band system inherently displays inversion symmetry. The Taylor expansion of the Bessel functions in the electric field also allows us to gauge this. Because they have a very particular form of expansion $J_{\alpha}\left(z\right)=\sum_{n=0}^{\infty}C_{n}z^{2n+\alpha}$, it is not hard to see from the coefficients in Eq. 14 that only odd powers of the electric field amplitude appear, which means that the current is inverted if the electric field is inverted.
@@
Notice that now, in the clean limit $\gamma\rightarrow 0$, the second and third terms of Eq. 13 do not decay and actually dominate the response since the first term vanishes in this limit. It is evident from Eq. 15 that one can generate even harmonics for $\phi\neq 0$. Although the emergence of even harmonics is usually thought to be prohibited in centrosymmetric systems, this effect is solely due to the sudden switching of the electric field. If the switching is rather done adiabatically, the even harmonic contributions vanish (see Appendix VI.1). By introducing dissipation, the system eventually loses memory of the electric field’s initial condition and the even harmonics vanish. Furthermore, for $\phi=0$, the response is constituted solely of odd harmonics of amplitude $\chi_{2j+1}^{\gamma=0,\mathbf{l},1}=2J_{2j+1}^{\mathbf{l}}$ , as reported in Ref. [ghimire_observation_2011].
@@ II.3 Constant field
which as we will see is a generalization of the Esaki-Tsu relation [esaki_superlattice_1970]. The physical meaning of this is easier to unveil when considering a one dimensional tight binding chain with nearest neighbour hoppings. In that case we only have contribution from neighbours located at $l=1$ and the frequency $\omega_{B}^{\mathbf{l}}=\frac{e}{\hbar}E$ reduces to the Bloch frequency. The transient time is therefore composed of decaying Bloch Oscillations (BOs). These decaying BOs have also been predicted to occur in two-terminal mesoscopic setups subject to strong electric fields [pinho_bloch_2023], where the BOs decay is caused by the coupling of the mesoscopic system to external leads. If no dissipation was present, $\gamma=0$, then Eq. 17 would give a total current of $J\left(t\right)=-2t_{1}C\sin\left(\omega_{B}t\right)$, which are the expected persistent BOs predicted for the one-dimensional tight binding chain [Bloch29, Zener1934]. The constant, long-time current for the chain is $J^{\text{s}}=-2Ct_{1}\frac{E/\gamma}{\left(\frac{\hbar}{e}\right)^{2}+\left(E/\gamma\right)^{2}},$ which is the Esaki-Tsu relation that has been already obtained for one-dimensional nearest-neighbour tight binding lattices [kolovskyAtomicCurrentOptical2008, han_solution_2013, minot_quantum_2004]. The result in Eq. 18 is the generalization of the Esaki-Tsu formula for an arbitrary single-band tight-binding system.
@@ III Results
The electric field is considered to lay along the direction of the first neighbor which lies in the x axis and we also measure the current along this direction. The presence of an additional hopping gives rise to competing scales that generate interesting phenomena [ghimire_observation_2011].
@@ III.1 Monochromatic field
In Fig. 2 we plot the ratio between the odd harmonic Fourier amplitudes and the first-harmonic amplitude $\left|\Lambda_{j}^{\mathbf{l}}\right|/\left|\Lambda_{1}^{\mathbf{l}}\right|$ of a single neighbor for a set of dissipation values $\gamma/\Omega$ and for $\omega_{B}^{\mathbf{l}}/\Omega=3,12.2$ and $32$. A first thing to notice is how the amplitudes start dying out exponentially for harmonics of order superior than $\omega_{B}^{\mathbf{l}}$. By analysing Eq.(14), one can understand this behaviour as originating from the properties of Bessel functions. At a fixed argument $z$, as a function of the order $n$, $J_{n}\left(z\right)$ decays exponentially as $J_{n}\left(z\right)\sim e^{-n\ln\left(z^{-1}\right)}$ for large order $n$ [abramowitzHandbookMathematicalFunctions1964]. Another interesting result is how for strong dissipation the higher harmonics are suppressed, when compared to the first harmonic. This result is in agreement with results from Ref. [brown_real-space_2024], where it is interpreted as being a consequence of the suppression of a particle’s longer trajectories in real-space via dissipation, which are responsible for hi˛gh frequency contributions. However, our results indicate that the suppression of higher harmonics with dissipation is not monotonous with increased dissipation. As shown in Fig. (2)b), intermediate harmonics can have their amplitude increased with dissipation.
@@
are much smaller than the odd ones. However, for $\phi\neq 0$ the even harmonics amplitude approaches a finite value in the limit $\gamma/\Omega\ll 1$.In Fig. 4, we plot the transient and long times Fourier spectra for a system with $\omega_{B}^{\mathbf{l}}/\Omega=2.5$, $\Omega=0.13$, $\gamma=0.004$ and $\phi=0$ (Fig. 4b) and $\phi=\pi/2$ (Fig. 4c). The difference between the transient and long times spectra is the aquisition time of the signal, which is kept close or far away from the regime where the transient current dominates, respectively. A small second harmonic is visible in the transient spectrum for $\phi=0$, with amplitude of roughly two orders of magnitude smaller than that of the first harmonic. For $\phi=\pi/2$, however, the second harmonic is comparable to the first in the transient spectrum and a small DC current is also present. As the aquisition time increases, the oscillatory long time state has a more significant contribution and the even harmonics decrease with respect to the first harmonic in both plots.
@@
Figure 4: a) Time-evolution of current for $\omega_{B}^{\mathbf{l}}/\Omega=1.54$, $\Omega=0.15$ and $\gamma=0.003$. The black dashed line delimits the aquisition time of for the transient regime spectrum. b) and c) Transient and long times Fourier spectra for $\omega_{B}^{\mathbf{l}}/\Omega=2.5$, $\Omega=0.13$ and $\gamma=0.004$ in logarithmic scale for electric field phases b) $\phi=0$ and c) $\phi=\pi/2$. The aquisition times for the transient and long times spectra were 20002000 and $42400$ for both plots. A small second harmonic is visible in the transient spectrum in b), with amplitude of roughly two orders of magnitude smaller than that of the first harmonic. . In c), however, the second harmonic is comparable to the first harmonic in the transient spectrum and a small DC current is also present. As the aquisition time increases, the oscillatory long time state has a more significant contribution and the even harmonics decrease with respect to the first harmonic in both plots.
Although the emergence of even harmonics for centrossymetric systems is bound to exist solely on the transient regime, one can think of a setup aiming at consistently exciting the transient state, and therefore the even harmonics, by periodically pumping the system with optical pulses. The caveat for observing this phenomenon is that the switching of the electric field must be faster than its carrying period, as discussed in Appendix VI.1. One possible way of realizing this condition is by using a train of square-like pulses. The corresponding electric field can be written as
$E_{\text{pulses}}(t)$ $=\frac{E}{4}\sin(\Omega t)\sum_{n=1}^{N_{\text{pulses}}}\left[\operatorname{erf}\!\left(\frac{t-(s_{n}-w)}{\sigma}\right)+1\right]\times$ (20)
$\qquad\qquad\qquad\qquad\times\left[-\operatorname{erf}\!\left(\frac{t-(s_{n}+w)}{\sigma}\right)+1\right]$
Where $\text{erf}(x)$ is the error function, $s_{n}$ is the center of pulse $n$, $w$ is the pulses’ half-width and $\sigma$ controls hwo fast the pulse is switched on/off. We consider a train where the pulses are spaced equally, i.e, $s_{n}=s_{0}+n\Delta s$ and their spacing and width are multiples of the carrying period $T=2\pi/\Omega$ so as to respect the phase $\phi=\pi/2$ for each pulse. An example of the pulse train and its corresponding current response are shown in Fug. 5a) (in this case, the current was obtained by numerically integrating Eq. 12). The system is pumped out of equilibrium into the transient state by the first pulse during its width. After the pulse is switched off, the system is allowed to relax back into the initial state before being excited again by the upcoming pulse, allowing the excitation of the same transient state. Therefore, the pulses’ spacing, width and carrying frequency can be tuned for a given dissipation strength to allow the consistent emergence of strong even harmonics. Fig. 5b) depicts the Fourier spectrum (and its corresponding time profile in Fig. 5b0) ) of the current response for a given impulse, showcasing a second and fourth harmonics of the same orders of magnitude as the first and third, respectively.
Figure 5: a) An example of a train of square electric pulses (solid blue line) and its corresponding current response (solid red line) for $\omega_{B}^{\mathbf{l}}/\Omega=5$, $\Omega=0.05$ and $\gamma=0.002$. The pulse parameters are $\sigma=10$, $w=8\pi/\Omega$ and $\Delta s=34\pi/\Omega$. b) Fourier spectra for $\omega_{B}^{\mathbf{l}}/\Omega=2.5$, $\Omega=0.1$ and $\gamma=0.0005$ in logarithmic scale. The pulse train is composed of 12 square pulses with $\sigma=10$, $w=100\pi/\Omega$ and $\Delta s=240\pi/\Omega$. The corresponding time profile of the current and electric field are plotted in b0). Strong even harmonics are detected in the Fourier spectrum.